OpenStudy (anonymous):

Find indefinite intergral of x^3

zepdrix (zepdrix):

Are you familiar with the `Power Rule for Integrals`: \[\large \int\limits x^n\;dx \qquad=\qquad \frac{x^{n+1}}{n+1}\] ~Increase the exponent by 1, ~then divide by this new exponent. Example:\[\large \int\limits x^2\;dx \qquad=\qquad \frac{x^{2+1}}{2+1} \qquad=\qquad \frac{1}{3}x^3\] Do you understand how you could apply this to your problem? :o

OpenStudy (anonymous):

This works really well with simple functions like x^3